You are considering how to invest part of your retirement savings. You have decided to put $ 500 comma 000 into three​ stocks: 66 % of the money in GoldFinger​ (currently $ 27 ​/share), 19 % of the money in Moosehead​ (currently $ 83 ​/share), and the remainder in Venture Associates​ (currently $ 1 ​/share). Suppose GoldFinger stock goes up to $ 38 ​/share, Moosehead stock drops to $ 52 ​/share, and Venture Associates stock rises to $ 18 per share. (Round to the nearest​ dollar.)

a. What is the new value of the​ portfolio? (Round to the nearest​ dollar.)

b. What return did the portfolio​ earn?

c. If you​ don't buy or sell any shares after the price​ change, what are your new portfolio​ weights?
Answer #1
Shares inv. in Goldfinger = Amnt inv.*%age in Goldfinger/price of Goldfinger=500000*0.66/27=12222.2222
Shares inv. in Moosehead = Amnt inv.*%age in Moosehead/price of Moosehead=500000*0.19/83=1144.5783
Shares inv. in Venture ass. = Amnt inv.*%age in Venture ass./price of Venture ass.=500000*0.15/1=75000
A.
New portfolio value= Goldfinger new price*Goldfinger shares + Moosehead new price*Moosehead shares + Venture ass. new price*Venture ass. shares
=38*12222.2222+52*1144.5783+18*75000
=1873963
B.
Return % = New portfolio value/old portfolio value-1
=1873962.52/500000-1
=274.79%
C.
New weight of Goldfinger = New price*shares/new porfolio value
=38*12222.2222/1873962.52
=0.2478 = 24.78%
New weight of Moosehead = New price*shares/new porfolio value
=52*1144.5783/1873962.52
=0.0318 = 3.18%
New weight of Venture ass. = New price*shares/new porfolio value
=18*75000/1873962.52
=0.7204 = 72.04%